OpenStudy (anonymous):

write the equation of a line thats has a slope of -3 and passes through (4,5)

OpenStudy (anonymous):

y=-3x+5

OpenStudy (anonymous):

show work please and i have more

OpenStudy (accessdenied):

You have the slope and a point. Do you know the setup for point-slope form of a line?

OpenStudy (anonymous):

yes

OpenStudy (accessdenied):

So we have: \( y - y_1 = m(x - x_1) \) The point we are given is set in place of \((x_1, y_2)\) and the slope is set in place of m. Makes sense?

OpenStudy (accessdenied):

The point (4, 5) is not a y-intercept.

OpenStudy (anonymous):

yes but what then?

OpenStudy (accessdenied):

\(y - y_1 = m(x - x_1) \) \(y - 5 = -3 (x - 4) \) Technically, the question gives no goal about what form to leave it in. This would actually be sufficient as the equation describing the line.

OpenStudy (accessdenied):

5 is the y-value of the point (4, 5) But the point, (4, 5), is not on the y-axis. That is what we refer to by the y-intercept, is the intersection of the line and the y-axis.

OpenStudy (accessdenied):



OpenStudy (anonymous):

OH

OpenStudy (anonymous):

i have another

OpenStudy (accessdenied):

Go ahead. :)